OpenStudy (blackops2_143):

evaluate m - n when m = 3.4 and n = -8.7

OpenStudy (alexandervonhumboldt2):

substitute the values: m-n=3.4-(-8.7)=3.4+8.7=12.1

OpenStudy (trwatkins1):

12.1